The length of a rectangle is 6 feet less than 3 times its width. The area is 360 square feet. What is the length of the rectangle?
Answer by mananth(16946)

let width=x
length = 3x-6
..
Area of rectangle = L*W
..
x(3x-6)=360
3x^2-6x=360
3x^2-6x-360=0
x^2-2x-120=0
x^2-12x+10x-120=0
x(x-12)+10(x-12)=0
(x-12)(x+10)=0
x=12 the width
..
length = 3x-6
=3*12-6
=36-6
=30 feet